# Scalable Multi-Objective Reinforcement Learning with Fairness Guarantees using Lorenz Dominance

### v1

@@ Abstract
Multi-Objective Reinforcement Learning (MORL) aims to learn a set of policies that optimize trade-offs between multiple, often conflicting objectives. MORL is computationally more complex than single-objective RL, particularly as the number of objectives increases. Additionally, when objectives involve the preferences of agents or groups, ensuring fairness is socially desirable. This paper introduces a principled algorithm that incorporates fairness into MORL while improving scalability to many-objective problems. We propose using Lorenz dominance to identify policies with equitable reward distributions and introduce $\lambda$-Lorenz dominance to enable flexible fairness preferences. We release a new, large-scale real-world transport planning environment and demonstrate that our method encourages the discovery of fair policies, showing improved scalability in two large cities (Xi’an and Amsterdam). Our methods outperform common multi-objective approaches, particularly in high-dimensional objective spaces.
1 Introduction
@@
Reinforcement Learning (RL) is a powerful framework for sequential decision-making, in which agents learn to maximize long-term rewards by interacting with an environment (?). In typical RL applications, rewards are formalized by combining multiple criteria (or objectives) into a single scalar value, by weighting them (?). This approach, however, assumes prior knowledge of precise preferences over these objectives, which may not always be available in real-world scenarios. Additionally, real-world problems often involve multiple, often conflicting objectives. Defining a scalar reward function prior to training can bias the decision-making process, potentially overlooking valuable policies that vary primarily in how these objectives are weighted (?).
@@
Multi-Objective Reinforcement Learning (MORL) addresses this challenge by introducing a distinct reward function for each objective (?). This results in a set of candidate optimal policies, rather than a single solution, allowing decision-makers to select policies based on their private preferences. MORL has been successfully applied to various domains, including decision-making under unknown preferences (?, ?), human-value alignment (?, ?), robot locomotion (?), and multi-agent systems (?, ?, ?).
@@
Multi-policy methods in MORL address situations with unknown preference settings by assuming a monotonically increasing utility function and optimizing for all objectives simultaneously, approximating the Pareto front of optimal, efficient policies (?, ?, ?). However, even with this minimal assumption, the solution set scales exponentially with the number of objectives, a challenge that is particularly pronounced in many-objective optimization, where the problem involves many different objectives (?, ?). As a result, multi-policy methods often struggle to scale efficiently in such scenarios, underscoring the importance of further research in many-objective optimization within RL (?).
@@
Furthermore, learning the complete Pareto front of a problem is often unnecessary, as certain policies may be inherently undesirable (?). For example, in fairness-critical applications, some Pareto-non-dominated policies may lead to inequitable reward distributions between different objectives, which is particularly undesirable when objectives represent the utilities of different societal groups or people (?, ?). Although egalitarian approaches, such as maxmin or equal weighting can address this, they assume a predefined exact preference over the different objectives, often limiting flexibility and can sometimes result in inefficient solutions (?). This reveals a research gap in MORL: no current multi-policy method (a) provides fairness guarantees to the decision-maker, (b) allows control over the fairness constraint, and (c) scales effectively to problems with many objectives.
@@
In this paper, we propose leveraging Lorenz dominance to identify a subset of the Pareto front that ensures equitable reward distribution without predefined preferences. We extend this approach with $\lambda$-Lorenz dominance, which allows decision-makers to adjust the strictness of fairness constraints via a parameter $\lambda$. We formally demonstrate that $\lambda$-Lorenz dominance enables interpolation between Lorenz and Pareto dominance, offering decision-makers fine-grained control over fairness constraints. Furthermore, we introduce Lorenz Conditioned Networks (LCN), a novel algorithm designed to optimize for $\lambda$-Lorenz dominance.
@@
To support scalability in many-objective settings, we design a new large-scale, multi-objective environment for planning transport networks in real-world cities with a flexible number of objectives. We conducted experiments in the cities of Xi’an (China) and Amsterdam (Netherlands) and show that LCN generates fair policy sets in large objective spaces. As Lorenz-optimal sets are generally smaller than Pareto-optimal sets (?), LCN exhibits excellent scalability, especially for many-objective problems.
Alongside the paper, we release the code and data used to generate our results  [^1] Github repository: https://github.com/sias-uva/mo-transport-network-design.
2 Related Work
@@
Our work lies at the intersection of multi-policy methods for MORL (?) and algorithmic fairness in sequential decision-making problems (?).
2.1 Multi-Policy MORL
@@
Early multi-policy methods in reinforcement learning (RL) used Pareto Q-learning, which was mainly limited to small-scale environments (?, ?, ?). To address scalability, many approaches assume that decision makers have linear preferences, leading to a simpler solution set known as the convex coverage set (?, ?, ?). For example, GPI-LS, a state-of-the-art method against which we compare our method, decomposes the multi-objective problem into single-objective subproblems in which the reward function is a convex combination of the original vectorial reward and subsequently trains a neural network to approximate the optimal policies under different weights (?). Beyond linear preferences, the Iterated Pareto Referent Optimisation (IPRO) method takes a similar decomposition-based approach and provides strong theoretical guarantees but was shown to suffer when increasing the number of objectives (?). In contrast, Pareto Conditioned Networks (PCNs) do not decompose the problem, but instead train a return-conditioned policy(?, ?, ?). Similarly, PD-MORL approximates the Pareto front by uniformly sampling preferences across the preference space (?). We propose a method inspired by PCNs that focuses on fairness, avoiding the need to search the entire preference space. This allows for scalability to higher dimensions while learning a set of fair policies, and offers flexibility in setting fairness preferences.
2.2 Fairness in MORL
@@
Research on fairness in RL can be categorized along two main themes (?): fairness in domains where individuals belong to protected groups (societal bias) and fairness in resource allocation problems (non-societal bias). Our work aligns closely with the first theme, focusing on the fair distribution of benefits among different societal groups. Group fairness has been studied in RL before, specifically in multi-agent scenarios (?, ?), where agents learn individual policies. Although here we focus on single-agent RL, we assume that the policies learned by the agent will impact groups of individuals, possibly with conflicting preferences.
@@
Fairness in RL often involves balancing multiple objectives. Many studies in this area incorporate diverse objectives into a single fairness-based reward function. This is typically achieved through linear reward scalarization (?, ?, ?), generalization to nonlinear combinations of rewards, and welfare functions such as the Generalized Gini Index (?, ?, ?), as well as other reward-shaping mechanisms (?, ?, ?, ?). Alternatively, some methods modify the reward function during training to comply with fairness constraints (?). These approaches require encoding the desired fairness principles into the reward functions beforehand, requiring preferential information prior to training. Our approach does not make such assumptions.
@@
Our work is closely related to (?), who propose a formal MORL fairness framework that encodes six fairness notions as objectives. They train PCN to identify Pareto-optimal trade-offs between these fairness notions. Although our method can be applied within this framework, it differs in that it does not precalculate any specific fairness notion. Instead, it learns a set of non-dominated policies across all objectives, allowing the decision-maker to define their criteria post-training. While Lorenz-dominance has previously been used in multi-objective optimization methods (?, ?), this paper is the first attempt to use its properties to learn multiple policies in MORL.
3 Preliminaries
@@
In this section, we formally introduce multi-objective reinforcement learning and the fairness criterion we focus on.
3.1 Multi-Objective Reinforcement Learning
@@
In essence, ${\bm{v}}$ Pareto dominates ${\bm{v}}^{\prime}$ when it is at least equal in all objectives and better in at least one. For a set of vectors $D$, the Pareto front $\mathcal{F}\left(D\right)$ contains all vectors that are non-Pareto dominated.
3.2 Fairness in Many-Objective Reinforcement Learning
One common fairness approach in multi-objective reinforcement learning (MORL) is to treat all objectives as equally important, optimizing a single, equally weighted objective. However, this assumes absolute equality in reward distribution, which may be infeasible in certain problems, and yields a single policy, without offering options to the decision-maker. Another approach, inspired by Rawlsian justice theory and the maxmin principle, focuses on maximizing the minimum reward between objectives. However, this often results in solutions that are not fully efficient for all users (?).
To train a multi-policy algorithm with fair trade-offs, it is necessary to identify all optimal trade-offs that achieve a fair distribution of rewards. To achieve this, we use Lorenz dominance, a refinement of Pareto dominance that considers the distribution of values within a vector (?). This concept, traditionally used in economics to assess income inequality (?), is adapted here for fairness in MORL.
@@ 3.2 Fairness in Many-Objective Reinforcement Learning
where $v_{(1)}\leq v_{(2)}\leq...\leq v_{(d)}$ are the values of the vector ${\bm{v}}$, sorted in increasing order. A vector ${\bm{v}}$ Lorenz dominates a vector ${\bm{v}}^{\prime}$, when its Lorenz vector $L({\bm{v}})$ Pareto dominates the Lorenz vector $L({\bm{v}}^{\prime})$ (?).
Our fairness criterion seeks Pareto-optimal trade-offs that distribute rewards fairly across objectives, without encoding this preference explicitly into a scalarization function. This approach generates a set of trade-offs, allowing the decision-maker to select the most appropriate one after training. Lorenz-based fairness is grounded in the Pigou-Dalton transfer principle from economics (?, ?). For a vector ${\bm{v}}\in\mathbb{R}^{d}$ with $v_{i}>v_{j}$ for some $i,j$, a transfer of any $\epsilon$ ($0<\epsilon\leq v_{i}-v_{j}$) to create ${\bm{v}}’={\bm{v}}-\epsilon I_{v_{i}}+\epsilon I_{v_{j}}$, where $I_{v_{i}}$ and $I_{v_{j}}$ are indicator vectors for the $i$th and $j$th elements, respectively, results in a preferred distribution of rewards while preserving the total sum (?). Consider, for example, vector ${\bm{v}}=(8,0)$. If we perform a transfer of $\epsilon=4$, the new vector ${\bm{v}}^{\prime}=(4,4)$ is more desirable under Lorenz-fairness, since the reward is distributed equally between the two dimensions, while the total sum remains the same.
@@
In MORL, we define vectors ${\bm{v}}^{\pi},{\bm{v}}^{\pi^{\prime}}$ as the expected return of the policies $\pi,\pi^{\prime}$, across all objectives of the environment, respectively. We define fair policies as those that are non-Lorenz dominated. The set of non-dominated value vectors is called a Lorenz coverage set, which is usually (but not necessarily) significantly smaller than a Pareto coverage set (?). Our fairness approach satisfies the criteria outlined by ?. It is Lorenz-efficient as the learned policies are non-Lorenz-dominated; it is impartial, since it treats all objectives as equally important, and it is equitable, as Lorenz dominance satisfies the Pigou-Dalton principle (?). In Figure 1, we show the difference between Pareto and Lorenz dominance. The latter extends the are of undesired solutions, allowing for fewer non-dominated solutions and providing fairness guarantees in the two objectives.
@@
Figure 1: The Pareto and Lorenz-dominated areas of vector S. The Lorenz-dominated area includes the Pareto-dominated area, and is symmetric around the equality line, except for the symmetric vector S’ = (3, 2). This creates an expanded dominance, resulting in fewer acceptable trade-offs.
4 Flexible Fairness with $\lambda$-Lorenz Dominance

### v2

@@ Abstract.
Multi-Objective Reinforcement Learning (MORL) aims to learn a set of policies that optimize trade-offs between multiple, often conflicting objectives. MORL is computationally more complex than single-objective RL, particularly as the number of objectives increases. Additionally, when objectives involve the preferences of agents or groups, incorporating fairness becomes both important and socially desirable. This paper introduces a principled algorithm that incorporates fairness into MORL while improving scalability to many-objective problems. We propose using Lorenz dominance to identify policies with equitable reward distributions and introduce $\lambda$-Lorenz dominance to enable flexible fairness preferences. We release a new, large-scale real-world transport planning environment and demonstrate that our method encourages the discovery of fair policies, showing improved scalability in two large cities (Xi’an and Amsterdam). Our methods outperform common multi-objective approaches, particularly in high-dimensional objective spaces.
copyright: ccdoi: 10.1613/jair.1.19862journalvolume: 85article: 31publicationmonth: 3journalyear: 2026
1. Introduction
@@
Reinforcement Learning (RL) is a powerful framework for sequential decision-making, where agents learn to maximize long-term rewards by interacting with an environment (wang_deep_2020). In most RL applications, rewards are constructed by aggregating multiple criteria (or objectives) into a single scalar value, typically via a weighted sum (hayes_practical_2022). However, this approach assumes prior knowledge of the precise preferences among objectives—a condition that rarely holds in real-world settings. Moreover, many real-world problems inherently involve multiple, often conflicting, objectives. Defining a scalar reward function before training can therefore introduce bias into the learning process, potentially excluding policies that differ primarily in their objective weightings (vamplew_scalar_2022).
@@
Multi-Objective Reinforcement Learning (MORL) addresses this challenge by defining a separate reward function for each objective (hayes_practical_2022). This yields a set of candidate optimal policies, rather than a single solution, that decision-makers select according to their preferences. MORL has been successfully applied in various domains, such as decision-making under unknown preferences (roijers_computing_2015; alegre_optimistic_2022), human-value alignment (peschl_moral_2021; rodriguez-soto_instilling_2022), robot locomotion (cao2021efficient), and multi-agent systems (radulescu_equilibria_2019; ropke_reinforcement_2023; vamplew_scalar_2022).
@@
Single-policy MORL learns one policy based on predefined knowledge about decision-makers’ preferences. However, such preferences are not always known at training time. Multi-policy methods in MORL handle unknown preferences by assuming a monotonically increasing utility function and optimizing all objectives simultaneously, approximating the Pareto front of optimal policies (mannion_multi-objective_2021; hayes_practical_2022; reymond_pareto_2022). Multi-policy methods face scalability challenges as the solution set can scale exponentially with the number of objectives. This issue becomes particularly severe in many-objective optimization, where the number of objectives is large (nguyen_multi-objective_2020; perny_approximation_2013). Consequently, multi-policy methods often struggle to scale efficiently in these scenarios, highlighting the need for further research in many-objective RL (hayes_practical_2022).
@@
Learning the entire Pareto front is often unnecessary, since some policies may be inherently undesirable (osika2023what). For example, in fairness-critical applications, some Pareto-non-dominated policies may result in unequal reward distributions across objectives. This is especially problematic when objectives represent the utilities of different societal groups (pmlr-jabbari-fairness; cimpean_multi-objective_2023). Although egalitarian approaches such as maxmin or equal weighting can address this issue, they assume a predefined, exact preference over objectives, often limiting flexibility and sometimes yielding inefficient results (siddique2020learning). This reveals a research gap in MORL: no current multi-policy method (a) guarantees fairness to the decision-maker, (b) allows control over fairness constraints, and (c) scales effectively to many-objective problems.
@@
In this paper, we propose using Lorenz dominance to identify a subset of the Pareto front that ensures equitable reward distribution, without requiring predefined preferences. We extend this approach with $\lambda$-Lorenz dominance, enabling decision-makers to adjust the strictness of fairness constraints through a parameter $\lambda$. We formally show that $\lambda$-Lorenz dominance interpolates between Lorenz and Pareto dominance, providing decision-makers with fine-grained control over the degree of fairness. We also introduce Lorenz Conditioned Networks (LCN), a novel algorithm for optimizing $\lambda$-Lorenz dominance.
@@
To support scalability in many-objective settings, we develop a new large-scale, multi-objective environment for planning transport networks in real-world cities with a flexible number of objectives. We conduct experiments in the cities of Xi’an (China) and Amsterdam (Netherlands) and show that LCN generates fair policy sets in large objective spaces. Since Lorenz-optimal sets are typically smaller than Pareto-optimal sets (perny_approximation_2013), LCN scales effectively, particularly in many-objective problems. We release the code and data used to generate our results alongside this paper  [^1] GitHub repository: https://github.com/sias-uva/mo-transport-network-design.
2. Related Work
@@
Our work intersects multi-policy methods for MORL (hayes_practical_2022) and algorithmic fairness in sequential decision-making problems (gajane_survey_2022).
2.1. Multi-Policy MORL
@@
Early multi-policy methods in RL, such as Pareto Q-learning, were limited to small-scale environments (moffaert_multi-objective_2014; ruiz-montiel_temporal_2017; parisi2016multi). To improve scalability, many approaches assume linear decision-maker preferences, resulting in a simpler solution set called the convex coverage set (roijers_computing_2015; abels_dynamic_2019; felten_decomposition). For example, GPI-LS—a state-of-the-art method and our baseline—decomposes the multi-objective problem into single-objective subproblems. Each subproblem uses a reward function that is a convex combination of the original vectorial reward. It then trains a neural network to approximate optimal policies for different weights (alegre_gpi). Beyond linear preferences, the Iterated Pareto Referent Optimization (IPRO) method uses a similar decomposition-based approach and provides strong theoretical guarantees. However, its performance degrades as the number of objectives increases (ropke2024divide). In contrast, Pareto Conditioned Networks (PCNs) do not decompose the problem, but instead train a return-conditioned policy, (reymond_pareto_2022; reymond_exploring_2022; delgrange_wae-pcn_2023). PCNs have been applied in various domains, including water management (osika2025multi), pandemic intervention policies (chen2025learning), autonomous cyber defence (o2025multi) and battery control (huoptimize). Similar to PCN, PD-MORL approximates the Pareto front by uniformly sampling preferences across the preference space (basaklar2022pd), while C-MORL (liu2024c) bridges constrained policy optimization and MORL to efficiently discover the Pareto front through parallel policy training. We propose a method inspired by PCNs that focuses on fairness, avoiding the need to search the entire preference space. This enables scalability to higher dimensions, learning a set of fair policies and offering flexibility in setting the degree of fairness preference.
2.2. Fairness in MORL
@@
Research on fairness in RL can be categorized along two main themes (gajane_survey_2022): fairness in domains where individuals belong to protected groups (societal bias) and fairness in resource allocation problems (non-societal bias). Our work aligns closely with the first theme, focusing on the fair distribution of benefits among different societal groups. Group fairness has been studied in RL before, specifically in multi-agent scenarios (SatijaLPP23; ju2023achieving), where agents learn individual policies. While we focus on single-agent RL, we assume that the agent’s policies will affect groups of individuals, who may have conflicting preferences
@@
Achieving fairness in RL often requires balancing multiple objectives. Many studies in this area incorporate diverse objectives into a single fairness-based reward function. This is typically achieved through linear reward scalarization (rodriguez-soto_multi-objective_2021; chen_same-day_2023; blandin2024group), nonlinear reward combinations, and welfare functions (e.g. the Generalized Gini Index) (siddique2020learning; hu_towards_2023; fan_welfare_2023), and other reward-shaping mechanisms (zimmer_learning_2021; yu_policy_2022; mandal_socially_2023; kumar_fairness_2023). Alternatively, some methods adapt the reward function during training to satisfy fairness constraints (chen_bringing_2021). These approaches require encoding fairness principles into the reward functions a priori, requiring preference information before training. Our method avoids these assumptions.
@@
Our work is closely related to (cimpean_multi-objective_2023), which proposes a formal MORL fairness framework that encodes six fairness notions as objectives. The authors use PCNs to identify Pareto-optimal trade-offs among these fairness notions. While our method can be used within this framework, it differs by not predefining any specific fairness notion. Instead, it learns a set of non-dominated policies across all objectives, allowing the decision-maker to define their fairness criteria after training and select a policy accordingly.
Our method relies on Lorenz dominance, a criterion that favors policies with balanced reward distributions. Lorenz dominance has previously been used in multi-objective optimization methods (CHABANE20191; fasihi_bi-objective_2023; bederina2024generating), but its application in MORL has been limited. (perny_approximation_2013) first introduced the Lorenz criterion in Multi-Objective Markov Decision Processes (MOMDPs), providing much of the theoretical foundation we rely on this work. However, their experiments were limited to small-scale, randomly generated MOMDPs. Since then, some works inspired by the Lorenz curve have emerged; for example, (siddique2020learning) uses the Generalized Gini Function to create a weighted sum in the Lorenz space for single-objective RL. Building on the framework of (perny_approximation_2013), we train a neural network to learn the full Lorenz front, enabling flexible degrees of fairness for the decision-maker and demonstrating scalability to significantly larger and more realistic environments.
3. Preliminaries
@@
In this section, we formally introduce Multi-Objective Reinforcement Learning (MORL) and Lorenz dominance.
3.1. Multi-Objective Reinforcement Learning
@@
With vector-based rewards, there is generally no single optimal policy in the usual sense of scalar rewards (e.g. policy maximizing reward). Instead, we learn a set of optimal policies using a dominance criterion. In MORL, Pareto dominance is commonly applied, resulting in a solution set called the Pareto front.
@@
In essence, $\bm{v}$} Pareto dominates $\bm{v}$$^{\prime}$ when it is at least as good in all objectives and strictly better in at least one. For a set of vectors $D$, the Pareto front $\mathcal{F}\left(D\right)$ contains all non-Pareto-dominated vectors.
Definition 0.
(Pareto front) Given a set of vectors $D\subseteq\mathbb{R}^{d}$, the Pareto front $\mathcal{F}\left(D\right)$ is the subset of vectors in $D$ that are not Pareto-dominated by any other vector in $D$. Formally,
$\mathcal{F}\left(D\right)=\{{\bm{v}}\in D\mid\nexists\;{\bm{v}}^{\prime}\in D\text{ such that }{\bm{v}}^{\prime}\succ_{P}{\bm{v}}\}.$
3.2. Fairness in Many-Objective Reinforcement Learning
One common fairness approach in Multi-Objective Reinforcement Learning (MORL) is to treat all objectives as equally important, optimizing a single, equally weighted objective. However, this assumes absolute equality in reward distribution, which may be infeasible in certain problems, and yields a single policy, without offering options to the decision-maker. Another approach, inspired by Rawlsian justice theory and the maxmin principle, focuses on maximizing the minimum reward between objectives. However, this often results in solutions that are not efficient for all users (siddique2020learning).
To train a multi-policy algorithm with fair trade-offs, it is necessary to identify all optimal trade-offs that achieve a fair distribution of rewards. To achieve this, we use Lorenz dominance, a refinement of Pareto dominance that considers the distribution of values within a vector (perny_approximation_2013). This concept, traditionally used in economics to assess income inequality (ranking_income), is adapted here for fairness in MORL.
@@ 3.2. Fairness in Many-Objective Reinforcement Learning
where $v_{(1)}\leq v_{(2)}\leq\cdots\leq v_{(d)}$ are the values of the vector $\bm{v}$}, sorted in increasing order. A vector $\bm{v}$} Lorenz dominates a vector $\bm{v}$$^{\prime}$, when its Lorenz vector $L({\bm{v}})$ Pareto dominates the Lorenz vector $L({\bm{v}}^{\prime})$ (perny_approximation_2013). We use ${\bm{v}}\succ_{L}{\bm{v}}^{\prime}$ to denote that $\bm{v}$} Lorenz dominates $\bm{v}$$^{\prime}$.
For a set of vectors $D$, the Lorenz front $\mathcal{L}\left(D\right)$ contains all vectors that are non-Lorenz-dominated.
Definition 0.
(Lorenz front) Given a set of vectors $D\subseteq\mathbb{R}^{d}$, the Lorenz front $\mathcal{L}\left(D\right)$ is the subset of vectors in D that are not Lorenz-dominated by any other vector in D. Formally,
$\mathcal{L}\left(D\right)=\{{\bm{v}}\in D\mid\nexists\;{\bm{v}}^{\prime}\in D\text{ such that }L({\bm{v}}^{\prime})\succ_{P}L({\bm{v}})\},$
where $L({\bm{v}})$ denotes the Lorenz vector of $\bm{v}$}, and $\succ_{P}$ is the Pareto dominance relation.
Our approach builds on the Pigou-Dalton transfer principle from economics (perny_approximation_2013; adler_pigou-dalton_2013). This principle states that a redistribution of value from a better-off to a worse-off component improves fairness, as long as it does not reverse their ranking. Formally, given a reward vector ${\bm{v}}\in\mathbb{R}^{d}$ with two components $v_{i}>v_{j}$ for some indices $i$ and $j$, transferring a small amount $\epsilon$ ($0<\epsilon\leq v_{i}-v_{j}$) from $v_{i}$ to $v_{j}$ yields a new vector ${\bm{v}}^{\prime}={\bm{v}}-\epsilon I_{i}+\epsilon I_{j}$, where $I_{i}\in\mathbb{R}^{d}$ is an indicator vector with a 1 in the $i$th position and 0 elsewhere (and similarly for $I_{j}$). This transformation preserves the total reward and ranking, but results in a more equitable distribution, and is therefore called a Pigou-Dalton transfer (perny_approximation_2013).
Lorenz-based fairness evaluates policies based on how equitably rewards are distributed across objectives. A solution is considered fairer if it can be obtained via a sequence of Pigou-Dalton transfers from another, implying that it Lorenz-dominates the other. For example, consider ${\bm{v}}=(8,0)$. A transfer of $\epsilon=3$ yields ${\bm{v}}^{\prime}=(5,3)$. Even though the total reward remains 8, and the rank between entries is preserved, $\bm{v}$$^{\prime}$ is considered fairer under Lorenz dominance.
@@
In MORL, we define vectors ${\bm{v}}^{\pi},{\bm{v}}^{\pi^{\prime}}$ as the expected return of the policies $\pi,\pi^{\prime}$, across all objectives of the environment, respectively. We define fair policies as those that are non-Lorenz-dominated. The set of non-dominated value vectors is called a Lorenz coverage set, which is usually (but not necessarily) significantly smaller than a Pareto coverage set (perny_approximation_2013). Our fairness approach satisfies the criteria outlined in (siddique2020learning). It is Lorenz-efficient as the learned policies are non-Lorenz-dominated; it is impartial, since it treats all objectives as equally important, and it is equitable, as Lorenz dominance satisfies the Pigou-Dalton principle (perny_approximation_2013). In Figure 1, we show the difference between Pareto and Lorenz dominance. The latter extends the area of undesired solutions, allowing for fewer non-dominated solutions and providing fairness guarantees in the two objectives.
[…]
@@
Figure 1. The Pareto and Lorenz-dominated areas of vector $\bm{v}$}. The Lorenz-dominated area includes the Pareto-dominated area, and is symmetric around the equality line, except for the symmetric vector ${\bm{v}}^{\prime}=(3,2)$. This creates an expanded dominance, resulting in fewer acceptable trade-offs.
4. Flexible Fairness with $\lambda$-Lorenz Dominance